# AudioStory: Generating Long-Form Narrative Audio with Large Language Models

### v1

@@ 1 Introduction
Figure 1: AudioStory decomposes multimodal instructions into a sequence of coherent audio segments, capturing scene transitions, emotional tone, and segment timestamps. Unlike prior T5-based diffusion models, which struggle with complex queries, AudioStory empowers LLMs with high-level planning ability for instruction-followed and consistent long audio generation.
@@ 2 Related Works
Any-to-any multimodal LLMs. Within the rapidly evolving field of multimodal learning, any-to-any generation across vision, language, and audio modalities represents a significant frontier jin2023unified ; tang2023codi2 ; wu2024nextgpt ; zhan2024anygpt ; lai2024spider ; ge2023seed-llama ; wu2024vila-u ; xie2024show-o ; team2024chameleon . This paradigm aims for models capable of accepting arbitrary input modalities and generating outputs in any desired modality. Pioneering efforts include CoDi tang2023codi and CoDi-2 tang2023codi2 , which leveraged composable diffusion for diverse modality handling. Spider lai2024spider further extended these capabilities by enabling the generation of multiple modalities in a single response. NExT-GPT wu2024nextgpt demonstrated the efficacy of lightweight alignment for adapting LLMs to multimodal tasks, while AnyGPT zhan2024anygpt showcased the potential of discrete sequence modeling for unified multimodal processing. Unified-IO2 lu2023unifiedio2 highlighted the impact of scale and unified architectures in achieving state-of-the-art performance across a broad spectrum of modalities and tasks. Despite these advancements, current methods exhibit limitations in long-context generation with complex instructions: First, they primarily focus on speech generation and simple caption-to-music or caption-to-sound tasks, struggling to comprehend general and intricate human instructions beyond basic caption; Second, their audio generation is typically limited to single, short segments, hindering the generation of longer audio sequences.
@@ 6 Extended Applications
While previous experiments have focused on text-based instructions, we now extend the application of AudioStory to a more practical scenario: video dubbing. This enhancement enables the model to thoroughly analyze video content, reason about the sequence of events and their corresponding timestamps, and generate synchronized audio. An initial approach is to employ Gemini-2.5-pro to generate a one-sentence caption summarizing the entire video, followed by instruction-based audio generation, as illustrated in Fig. 4. However, this method is not conducive to producing audio that aligns closely with the visual content. Furthermore, the model is designed to accept both video data and instructions as input. The LLM performs reasoning on the video and produces bridging tokens. During the reasoning phase, the LLM first understands the overall content of the video, then sequentially breaks it down into events based on their temporal order. It infers the specific visual details and corresponding audio information for each event. Technically, we replace the LLM with a pretrained video MLLM (i.e., Qwen2.5-VL bai2025qwen2.5vl ) and jointly train the LLM and audio generator using LoRA tuning. The training data is from the animated sound partition of AudioStory-10k. We provide the video dubbing results in Fig. 5.
@@
Given an audio segment and an instruction, our model performs audio continuation. AudioStory first reasons about the content of the subsequent audio to be generated, then proceeds with segment-by-segment generation. The concatenated results are shown in Fig. 6.
@@ Appendix A Implementation Details
Appendix B Qualitative Cases
For all cases, we separately generate multiple single audio clips and concatenate them to constitute the final long-form audio.
Instructional long-form audio generation and continuation.
First, we present more cases for long-form audio generation. Our model could automatically derive the duration of each audio segment to be generated, as shown in Fig. 7, Fig. 8 and Fig. 9. One could observe that AudioStory could accurately determine the number of events based on the instruction and provide precise descriptions for each audio segment, including both the events themselves and their associated sound effects. Finally, AudioStory can precisely reason out the caption for each segment and generate the corresponding audio clips accordingly.
@@ Appendix B Qualitative Cases
Figure 9: Audio continuation case #2.
The case of Tom&Jerry.
Given either an instruction or a video, our model can generate sound that not only aligns with the musical style of the Tom&Jerry cartoon but also vividly reflects the characters’ emotions and storyline, exhibiting strong generalization capabilities.
Video dubbing.
[…]

### v2

@@ 1 Introduction
Figure 1: AudioStory effectively follows multimodal instructions, decomposing them into a sequence of coherent audio segments, capturing scene transitions, emotional tone, and segment timestamps. Unlike prior T5-based diffusion models, which struggle with complex queries, AudioStory empowers LLMs with high-level planning ability for instruction-followed and consistent long audio generation.
@@ 2 Related Works
Any-to-any multimodal LLMs. Within the rapidly evolving field of multimodal learning, any-to-any generation across vision, language, and audio modalities represents a significant frontier jin2023unified ; tang2023codi2 ; wu2024nextgpt ; zhan2024anygpt ; lai2024spider ; ge2023seed-llama ; wu2024vila-u ; xie2024show-o ; team2024chameleon ; lin2025toklip ; ma2025genhancer ; guo2025aligned . This paradigm aims for models capable of accepting arbitrary input modalities and generating outputs in any desired modality. Pioneering efforts include CoDi tang2023codi and CoDi-2 tang2023codi2 , which leveraged composable diffusion for diverse modality handling. Spider lai2024spider further extended these capabilities by enabling the generation of multiple modalities in a single response. NExT-GPT wu2024nextgpt demonstrated the efficacy of lightweight alignment for adapting LLMs to multimodal tasks, while AnyGPT zhan2024anygpt showcased the potential of discrete sequence modeling for unified multimodal processing. Unified-IO2 lu2023unifiedio2 highlighted the impact of scale and unified architectures in achieving state-of-the-art performance across a broad spectrum of modalities and tasks. Despite these advancements, current methods exhibit limitations in long-context generation with complex instructions: First, they primarily focus on speech generation and simple caption-to-music or caption-to-sound tasks, struggling to comprehend general and intricate human instructions beyond basic caption; Second, their audio generation is typically limited to single, short segments, hindering the generation of longer audio sequences.
@@ 6 Extended Applications
Figure 7: Qualitative cases of audio continuation.
We present AudioStory’s extended applications, including instruction-conditioned narrative audio generation, audio continuation, and video dubbing. Specifically, AudioStory generates long-form narrative audio from diverse multimodal instructions, remains closely aligned with input video and directives, and exhibits robust instruction-following.
@@
While prior experiments focused on text-based instructions, we extend AudioStory to a more practical scenario: video dubbing. This enhancement lets the model thoroughly analyze video content, reason about event sequences with timestamps, and generate synchronized audio. An initial approach uses Gemini-2.5-pro to generate a video summary caption, then performs instruction-based long-form narrative audio generation for the given video, as shown in Fig. 4.
However, this method poorly aligns audio with visual content. Therefore, we further modified the model to accept both video data and instructions as input. Specifically, the LLM reasons about the video to generate bridging tokens: it first comprehends the video’s overall content, then sequentially decomposes it into temporally ordered events, inferring each event’s visual details and corresponding audio information. Technically, we replace the LLM with a pretrained video MLLM (i.e., Qwen2.5-VL bai2025qwen2.5vl ) and jointly train the LLM and audio generator using LoRA tuning. The training data is from the animated sound partition of AudioStory-10k. We provide the video dubbing results in Fig. 5.
@@
Given an audio segment and an instruction, our model performs audio continuation. AudioStory first reasons about the content of the subsequent audio to be generated, then proceeds with segment-by-segment generation. The cases are shown in Fig. 6 and Fig. 7.
@@ Appendix A Implementation Details
Appendix B More Qualitative Cases
we present more cases for long-form audio generation. Our model could automatically derive the duration of each audio segment to be generated, as shown in Fig. 8. It can be observed that AudioStory could accurately determine the number of events based on the instruction and provide precise descriptions for each audio segment, including both the events themselves and their associated sound effects. Finally, AudioStory can precisely reason out the caption for each segment and generate the corresponding audio clips accordingly.
@@ Appendix B More Qualitative Cases
Figure 10: Long-form audio generation case #3.